OpenStudy (anonymous):

Solve this equation by completing the square. 3x2 - 4x = -2

OpenStudy (anonymous):

@CliffSedge ... can you help? I have to sign off... thanks!

OpenStudy (anonymous):

Factor out the 3 to make the left side x^2-4x/3 to start.

OpenStudy (anonymous):

Then find half of -4/3, square it, multiply it by 3 and add it to both sides. 3(x^2-4x/3) + ___ = -2 + ___

OpenStudy (anonymous):

i don't know i have no idea

OpenStudy (anonymous):

You don't know how to square a fraction? Or you don't know how I got to this step?

OpenStudy (anonymous):

yeah i dont know how i got this step

OpenStudy (anonymous):

-8/3

OpenStudy (anonymous):

Alright, let's back up and look at what we're starting with. \[\large 3x^2-4x=-2\] There are a couple ways to complete the square, but I prefer to have the leading coefficient be 1, so I'm going to divide everything by 3. \[\large \rightarrow x^2-\frac{4}{3}x=-\frac{2}{3}\]

OpenStudy (anonymous):

for the blank i have -8/3

OpenStudy (anonymous):

3(x^2-4x/3) +-8/3= -2 + /8/3 its right

OpenStudy (anonymous):

Before we go on, a little background concept on what 'completing the square' means.